A person with a mass of 40kg runs up a flight of stairs in 12s. The flight of stairs has 20 steps and the height
of each step is 20cm.
a) How much work is done in climbing the stairs?
b) What is the power output of the person running up the stairs?
1
Expert's answer
2018-06-01T10:05:08-0400
W=mgh=40*9.81*20*0.2=1569.6 J – total work.
P=W/t=1569.6/12=130.8 w – power output.